### v1

@@ 2. Issue-2
    The LLM does not have enough context: Figure 12 shows an example where the underlying LLM was not able to add an additional API call (settings.reload()) required for correct migration. This is required because the target library, dynaconf, loads the environment variables when the library is imported for the first time, and a reload is required if the environment variables are changed. The original code uses python-dotenv to load the environment variables, which reads the environment variables at runtime. Notice how the test, which is in another file thus does not get passed to the LLM, updates the environment variables before running the tests, which is after the libraries are imported, causing the tests to fail without the settings.reload() call in the migrated code. If the LLM was given this test code that affects the migration-related code, it may have been able to understand that the reload is required.
    Original code
    1try:
    2 d.update(toml.loads(r))
    3except TypeError:
    4 try:
    5 d.update(toml.load(r))
    6 except TypeError:
    7 d.update(r)
    PyMigTool migration
    1try:
    2 d.update(tomli.loads(r))
    3except TypeError:
    4 try:
    5 d.update(tomli.load(r))
    6 except TypeError:
    7 d.update(r)
    Manual fix
    1try:
    2 d.update(tomli.loads(r))
    3except TypeError:
    4 try:
    5 d.update(tomli.load(r))
    6 except AttributeError:
    7 d.update(r)
    Figure 13: Example of manually replacing exception type in a toml→tomli migration.
  3. Issue-3
    The LLM does not always adopt to the target library’s behavior: Figure 13 shows an example where the LLM did not recognize that the target library raises a different exception in a scenario, and therefore it was unable to replace the exception type correctly.
  4. Issue-4
    The LLM skips part of the code: Despite explicitly mentioning in the prompt not to skip parts of the code (2), the LLM still skips part of the code in migrations, specially when the code pattern is repetitive or does not have source API usage. We observe the same behavior in Section 3, which inspired us to add the Merge skipped code post-processing step in PyMigTool to merge the skipped code with the migrated code (see paragraph 4.1.3.1). Because this step uses some heuristics that often cannot identify all skipped code, we still had to manually fix some skipped code.
4.3.2 Possible extensions of PyMigTool
The problems we discuss above can potentially be mitigated by the following extensions of PyMigTool:
  1. Extension-1
    Post-processing for reverse refactoring: We can add a post processing step that detects renamed public APIs, and renames them back to the original names, potentially solving Issue-1. Refactoring detection tools, such as RefactoringMiner RefactoringMiner2 , can be used to detect such refactoring changes, and refactoring tools, such as RefMerge refmerge , can be used to apply the changes.
  2. Extension-2
    Including all dependant code with the prompt: This will allow the LLMs to know whether and how the code to migrate is used in other parts of the project, including the test code, potentially resolving Issue-2. This approach, however, can be impractical for the files that have a large number of dependencies, due to the LLM’s context length limitations.
  3. Extension-3
    Reprompting with test error messages: When we manually fixed the migration in Figure 13 related to Issue-3, we were able to quickly identify the issue just by looking at the test error messages. This means, systematically reprompting PyMigTool with the test error messages may help it to understand the context better and produce a correct migration, similar to how SOAR ni2021soar uses error messages to improve their program synthesis process (see LABEL:sec:relatedwork:transformation). While this can potentially help PyMigTool to produce more correct migrations, the challenge is processing the error messages and generating a meaningful prompt for the LLM. Note that SOAR considers only two pairs of libraries, and therefore the error messages are not as diverse as in our case.
  4. Extension-4
    Designing a sophisticated merge skipped code algorithm: our merge skipped code algorithm only looks at the size of the skipped code and usages of source APIs (see paragraph 4.1.3.1). A more sophisticated algorithm that considers the context and the structure of the code, such as removed functions/classes, may potentially improve this step reducing the effect of Issue-4.
While the above steps can potentially mitigate the problems we identified, the positive side is that manually fixing the incomplete migrations was also simple, so further effort in more automation may not be necessary. For example, in Figure 13 (Issue-3), we were able to catch the issue easily just by looking at test error messages, and the fix was just replacing the exception type in the code. While reprompting with error messages may potentially fix the problem, it would also increase the cost of migration. The skipped code cases (Issue-4) are also fairly simple to resolve with modern IDEs. We use PyCharm’s change viewer during our manual fix, which allows us to easily identify the skipped code, and merge them with a click of mouse.
The case in Figure 12 (Issue-2) required us to look deep into the library documentation and was time consuming. However, the original developers are likely to be more familiar with the libraries and their behavior. Therefore, they may have been be able to fix the issues easily.
@@ 4. Extension-4
In some cases where migrations are initially marked as incorrect, the issues are related to configuration rather than code changes. For example, in a migration from chardet→cchardet, adjusting the Python version from 3.11 to 3.9 resolved the test failures. Similarly, we made a migration from flask→sanic successful by installing an additional dependency, sanic-ext, with no changes to the migrated code. The developers of the projects are likely to have been aware of such configuration requirements, therefore, they may be been able to fix the issues easily.
@@
Overall, we find that PyMigTool automatically completes migration for a large portion of cases (32% migrations) and does most of the complex migration work (86% changes) for the remaining migrations, often leaving simple edits for the developers to perform (14% changes). Further improvements in migration process may potentially help in reducing the manual effort. However, the manual effort seems fairly minimal in most cases, specially for the developers who are familiar with the libraries and the project. We stress that the author of this thesis, who has no prior knowledge of the projects, performed all the manual edits alone without help of any other developers. This indicates that the manual edit part can potentially be done by someone new to the project, or even be outsourced outside of the project team.
4.4 Threats to validity
This section discusses the potential threats to the validity of our study and the steps we take to mitigate them.
4.4.1 Internal validity
One author performs all manual fixes and their evaluation, which may introduce subjectivity or bias. To mitigate this, we follow a consistent process for manual fixes and rely on objective test outcomes as the primary measure of correctness.
Configuration and environment setup also pose internal threats. Some migrations initially fail due to environment or dependency issues rather than code errors. We address these by adjusting configurations when possible, but environment-specific factors, such as Python and library version, operating system, may still affect the results.
4.4.2 External validity
@@
We focus on Python projects that are well-tested and actively maintained, as selected through specific filters. This focus may limit the generalizability of our findings to projects with less test coverage, different maintenance status, or other programming languages. The set of library pairs and migration scenarios is also limited by the availability of analogous libraries and the manual validation process. As a result, the observed effectiveness of PyMigTool may not fully generalize to all migration contexts. That said, the size of the dataset (717 migrations) and the diversity of selected projects (175 repositories) provide a reasonable basis for evaluating the tool’s effectiveness in a variety of migration scenarios.
4.4.3 Construct validity
@@
A key threat to construct validity is the correctness of the automated and manual evaluation process. We use automated test suites to assess migration correctness, but these tests may not cover all possible behaviors or edge cases in the migrated projects. To minimize this, we select projects with at least 80% lines of code covered by tests, which helps ensure that the tests are comprehensive enough to provide a reliable measure of correctness.
@@
Our manual migration effort calculation metric (Equation 1) is based on lines of code changed, which may not capture all aspects of migration effort, such as time spent understanding or debugging the migration. We use it as a measurable objective proxy for the effort required.
@@ 5 Related Work
The majority of the existing library migration techniques focus on identifying API mappings between two analogous libraries. Some of these techniques use mine existing library migrations to identify the API mappings teyton2012mining ; teyton2013automatic ; alrubaye2019use ; alrubaye2019migrationminer ; alrubaye2020learning . Some techniques see the requirement of existing migrations as a limitation, and therefore use other techniques, such as machine learning alrubaye2018automating ; miningAnalogicalAPIs and natural language processing ni2021soar . These techniques only identify the API mappings, but do not perform client code transformation, which is a separate challenge.
@@
SOAR ni2021soar uses API documentation to identify API mappings, and then uses the mappings and unit tests to guide program synthesis for client code transformation. SOAR has several limitations that prevent it from being a general solution for library migration. SOAR is evaluated on only two pairs of deep learning libraries, focusing on their migration of neural network models. These APIs are usually called in a sequence, where the output of one call is used as input to the next call. The nature of these APIs allows SOAR to check the program state after each API call, which is not the case for most libraries. Additionally, given the limited number of libraries, their technique relies on the library-specific error message format to further guide the synthesis process, requiring extensive changes and technique adaptations for running SOAR on other libraries. SOAR supports one-to-many code changes for a small set of pre-determined APIs, whose information is hard-coded for the transformation process. SOAR also times out for 20% of the evaluated migrations, even for small code snippets, the largest being 183 lines of code.
@@
Zhou et al. hapim proposed HaPiM, which first trains an unsupervised machine learning model capable of API mapping, and then uses the API mappings to guide an LLM in code transformation. Similar to Chen et al. miningAnalogicalAPIs , HaPiM learns from API usage patterns of source and target libraries in their respective client projects, therefore does not require any existing migration examples. For the code transformation, they use the API mappings to prompt an LLM to replace source API calls with target API calls. Their approach outperforms MigrationMapper alrubaye2019use and GPT-3.5 Turbo on the BLEU bleu and CodeBLEU codebleu metrics evaluated on 5 Java library pairs. A significant limitation of HaPiM is that it works with small independent code snippets that use the source library APIs, and results in an equivalent snippet that uses the target library APIs. This is as opposed to real-life code where the source API usages are interleaved with other types of code, such as variable declarations, control statements, and other library usages. This means that to use HaPiM, developers need to first extract code snippets that contains only the source library API usages from the client code, then use HaPiM to replace them with target library APIs, and then re-integrate the transformed code back into the client code, which is not practical for large-scale library migrations.
@@
Nikolov et al. nikolov2025google describe Google’s experiences in using LLMs for several internal code transformation tasks, including Joda Time→Java Time migrations. They do not attempt to provide a general solution for library migration; rather, they use LLMs to perform the migration they need. Therefore, their prompts contains instructions and contextual information tailored specifically for Joda Time→Java Time migrations. Nonetheless, they show that using LLMs reduced migration time by at least 50% compared to manual migration.
@@
The above shows that there have been different approaches to automate the library migration process. However, these techniques come with one or more limitations that prevent them from being a general solution for library migration. Our tool, PyMigTool, resolves these limitations and can automate the entire library migration process for any arbitrary analogous library pair (Section 4). We evaluate PyMigTool on a large and diverse dataset of 717 migrations between 48 analogous library pairs, and show that PyMigTool can successfully perform most of the migration with no or minimal manual intervention from the developer.
@@
In this paper, we explore the capabilities of LLMs for automating Python library migration and build PyMigTool, a practical tool that combines LLMs with static and dynamic program analysis to automate the full migration process. In the first phase of the work, we conduct an empirical study using three LLMs on the PyMigBench dataset to understand how well LLMs can perform library migration. We find that the best performing LLM, 4o, correctly migrates 94% of the individual code changes. At the migration level, it perfectly migrates 57% of the migrations, and 94% of them have at least one correctly migrated code change. We also run unit tests for a subset of the migrations, and find that 64% of the migrations by 4o have the same sets of tests passing in the developer’s migration and the LLM’s migration. A much lower cost LLM Mini and the free LLama also perform relatively well, with both correctly migrating 89% of code changes. Overall, we find that LLMs can handle complex code changes, therefore, we conclude that LLMs are a promising approach for automating library migration. However, we also identify some challenges that need to be addressed to make LLM-based migration practical.
We use these insights to build PyMigTool, a CLI tool that automates the full library migration process in phase two. Our evaluation on 717 simulated migrations across 175 diverse projects and 48 library pairs shows that PyMigTool can fully automate a large portion of migrations (32%), with most remaining cases requiring only minor manual intervention (median 14% changes). The tool’s layered approach, including post-processing steps, improves LLM migration correctness (from 65% to 86%) and reduces developer effort. While PyMigTool is not able to handle every migration scenario perfectly, it lowers the barrier for large-scale library migration in real-world projects. Overall, PyMigTool represents a meaningful step toward making automated library migration accessible and effective for the software engineering community.
@@ Acknowledgements.
We thank Dr. Ildar Akhmetov, Northeastern University, for his contributions in to the early to the ideas of the work.

### v2

@@ 5.2 Evaluation metrics
  • The LLM does not have the dependent code: Figure 10 shows an example where the underlying LLM was not able to add an additional API call (settings.reload()) required for correct migration. This is required because the target library, dynaconf, loads the environment variables when the library is imported for the first time, and a reload is required if the environment variables are changed. The original code uses python-dotenv to load the environment variables, which reads the environment variables at runtime. Notice how the test, which is in another file that does not get passed to the LLM, updates the environment variables before running the tests. This updating happens after the library is imported, causing the tests to fail without the explicit settings.reload() call in the migrated code. If the LLM knew about the dependent code, it may be able to better understand the context of the migration.
@@
  • The LLM is unaware of configuration requirements: In some cases, the migration requires changes in the project configuration rather than code changes. For example, in a migration from chardet→cchardet, adjusting the Python version from 3.11 to 3.9 resolved the test failures. Similarly, we made a migration from flask→sanic successful by installing an additional dependency, sanic-ext, with no changes to the migrated code. The LLM is unaware of such configuration requirements, and therefore, it cannot perform such changes.
  • Often test expectations need to be updated: There are cases where the tests require changes due to the target library’s different behavior. For example, often the source library raises a generic exception, which the tests expect, while the target library raises a more specific exception. Another example is coloroma uses integers to represent colors (green is 32), while rich uses strings (green is "[green]").
  • MigrateLib does not handle mock migrations: Some tests mock the source library’s API calls in the tests, and these mocks need to be migrated to the target library’s API calls. This is something we did not consider when designing MigrateLib, and therefore, it does not handle such cases.
  • MigrateLib does not support multi-library migrations: We design MigrateLib to migrate from just one source library to just one target library, which may not always be sufficient. The project miguelgrinberg/microblog-api is such an example, which uses flask and some other libraries that depend on flask. To fully migrate to fastapi (or any other target library), we also need to migrate all these dependent libraries.
6 Discussion
@@
Overall, we find that MigrateLib automatically completes migration for a large portion of cases (32% migrations) and does most of the complex migration work (73% changes) for the remaining migrations, often leaving simple edits for the developers to perform (27% changes). Further improvements in migration process may potentially help in reducing the manual effort. However, the manual effort seems fairly minimal in most cases, specially for the developers who are familiar with the libraries and the project. We now discuss some potential improvement and future research and development directions for MigrateLib based on our findings.
The two post-processing steps currently included in MigrateLib shows that simple automated steps can significantly improve the correctness of migrations. However, we also discuss in Section 5.4 find that these implementations have limitations, therefore can be further improved. MigrateLib can also include a reverse-refactoring post-processing step to detect and revert unwanted refactorings. There are existing tools that can detect refactorings Tsantalis et al. (2022) and apply refactoring changes Ellis et al. (2022), which can be used to implement this step.
We provide the LLM with a fairly straight-forward context – the file that uses the source library APIs. This is often insufficient, and on the contrary, often too much if the file is large but only a small part of it needs migration. Identifying and sending more relevant part of the code may help the LLM to perform better migration. This however, comes with the cost of more pre-processing to identify the relevant code, as well as more post-processing to replace old code with the new code in the project.
Another potential improvement may be iteratively re-prompting the LLM with additional context, such as test error messages after running the tests on the migrated code. The challenge here is to identify meaningful error messages and generate a prompt that can help the LLM to understand the context better. Trying out different re-prompting strategies for this may be an interesting future research direction.
7 Threats to validity
7.1 Internal validity
There were several manual steps involved in our evaluations. Manual activities are inherently subjective and may contain errors. In Phase 1, we manually validated the correctness of code changes and labeled any non-migration related change as refactoring or non-refactoring. To minimize subjectivity, two authors independently perform these steps and resolved disagreements through discussion. In Phase 3, only one author performs all manual migrations. However, we rely on objective test outcomes as the primary measure of correctness.
The merge skipped code uses some thresholds to identify skipped code hunks, which we determine through experimentation. Different thresholds may yield different results. However, we believe our chosen thresholds strike a reasonable balance between adding back genuinely skipped code and avoiding incorrect additions. If the assumptions do not hold, the resulting migration may potentially fail, therefore, the results we have may be a lower bound of the actual migration correctness.
7.2 External validity
We use three off the shelf LLMs in Phase 1, and the best of them in Phase 3. Using an LLM specifically tuned for the library migration task could potentially yield better results. Therefore, our findings in this paper may be viewed as a lower bound of the performance of potentially more powerful and specialized models.
@@ 7.2 External validity
In Phase 3, we focus on Python projects that are well-tested and actively maintained, as selected through specific filters. This focus may limit the generalizability of our findings to projects with less test coverage, different maintenance status, or other programming languages. The set of library pairs and migration scenarios is also limited by the availability of analogous libraries and the manual validation process. As a result, the observed effectiveness of MigrateLib may not fully generalize to all migration contexts. That said, the size of the dataset (717 migrations) and the diversity of selected projects (175 repositories) provide a reasonable basis for evaluating the tool’s effectiveness in a variety of migration scenarios.
The libraries we consider in Phase 1 and Phase 3 are relatively popular and well-established. This familiarity with APIs and API usage is precisely our hope when using LLMs for migration. However, being familiar with the APIs alone does not always guarantee being able to map them to each other or apply the right transformations on all code bases, which is why our evaluation is necessary. For newer libraries, additional information, e.g., API documentation, may be needed to guide the LLM in the migration process. This can be an interesting avenue for future study, and thus improvement of MigrateLib.
7.3 Construct validity
In RQ-1.1, we analyze the code change categories based on Changedev, not on Changellm. The categories can differ if the LLM produces a different code change than the developer. PyMigBench already provides us with the Changedev categories; categorizing the LLM changes according to PyMigTax Islam et al. (2024) requires additional manual effort that is beyond the scope of this work. That said, our observation is that the code/API differences in most cases are minor, and would not yield in a different PyMigTax category.
@@ 7.3 Construct validity
We use automated test suites to assess migration correctness in Phase 3, but these tests may not cover all possible behaviors or edge cases in the migrated projects. To minimize this, we select projects with at least 80%lines of code covered by tests, which helps ensure that the tests are comprehensive enough to provide a reliable measure of correctness.
@@
We use migration lines of code (Equation 1) to estimate the amount of migration work, which may not capture all aspects of migration work, such as time spent understanding or debugging the migration. We choose to use this metric since it is a measurable objective proxy for the effort required.
@@ 8 Related Work
The majority of the existing library migration techniques focus on identifying API mappings between two analogous libraries. Some of these techniques mine existing library migrations to identify the API mappings Teyton et al. (2012, 2013); Alrubaye et al. (2019b, a, 2020). Some techniques see the requirement of existing migrations as a limitation, and therefore use other techniques, such as machine learning Alrubaye and Mkaouer (2018); Chen et al. (2019) and natural language processing Ni et al. (2021). These techniques only identify the API mappings, but do not perform client code transformation, which is a separate challenge.
@@
SOAR Ni et al. (2021) uses API documentation to identify API mappings, and then uses the mappings and unit tests to guide program synthesis for client code transformation. SOAR has several limitations that prevent it from being a general solution for library migration. SOAR is evaluated on only two pairs of deep learning libraries, focusing on their migration of neural network models. These APIs are usually called in a sequence, where the output of one call is used as input to the next call. The nature of these APIs allows SOAR to check the program state after each API call, which is not the case for most libraries. Additionally, given the limited number of libraries, their technique relies on the library-specific error message format to further guide the synthesis process, requiring extensive changes and technique adaptations for running SOAR on other libraries. SOAR supports one-to-many code changes for a small set of pre-determined APIs, whose information is hard-coded for the transformation process. SOAR also times out for 20% of the evaluated migrations, even for small code snippets, the largest being 183 lines of code.
@@
Zhou et al. Zhou et al. (2023b) proposed HaPiM, which first trains an unsupervised machine learning model capable of API mapping, and then uses the API mappings to guide an LLM in code transformation. Similar to Chen et al. Chen et al. (2019), HaPiM learns from API usage patterns of source and target libraries in their respective client projects, therefore does not require any existing migration examples. For the code transformation, they use the API mappings to prompt an LLM to replace source API calls with target API calls. Their approach outperforms MigrationMapper Alrubaye et al. (2019b) and GPT-3.5 Turbo on the BLEU Papineni et al. (2002) and CodeBLEU Ren et al. (2020) metrics evaluated on 5 Java library pairs. A significant limitation of HaPiM is that it works with small independent code snippets that use the source library APIs, and results in an equivalent snippet that uses the target library APIs. This is as opposed to real-life code where the source API usages are interleaved with other types of code, such as variable declarations, control statements, and other library usages. This means that to use HaPiM, developers need to first extract code snippets that contains only the source library API usages from the client code, then use HaPiM to replace them with target library APIs, and then re-integrate the transformed code back into the client code, which is not practical for large-scale library migrations.
@@
Nikolov et al. Nikolov et al. (2025) describe Google’s experiences in using LLMs for several internal code transformation tasks, including Joda Time→Java Time migrations. They do not attempt to provide a general solution for library migration; rather, they use LLMs to perform the migration they need. Therefore, their prompts contains instructions and contextual information tailored specifically for Joda Time→Java Time migrations. Nonetheless, they show that using LLMs reduced migration time by at least 50% compared to manual migration.
There are also a few studies that focus on migrating between different versions of the same library. The study of Nikolov et al. Nikolov et al. (2025) we describe above also includes JUnit 3→JUnit 4 migrations. Almeida et al. Almeida et al. (2024) evaluate the performance of ChatGPT for SQAlchemy 1→SQLAlchemy 2 migrations. While similar, migrating between different versions of the same library is different from migrating between different libraries, because the former can leverage the evolution or release history of the libraries to identify API changes.
@@
The above shows that there have been different approaches to automate the library migration process. However, these techniques come with one or more limitations that prevent them from being a general solution for library migration. Our tool, MigrateLib, resolves these limitations and can automate the entire library migration process for any arbitrary analogous library pair. We evaluate MigrateLib on a large and diverse dataset of 717 migrations between 48 analogous library pairs, and show that MigrateLib can successfully perform most of the migration with no or minimal manual intervention from the developer.
@@ 9 Conclusion
Our goal in this paper is to fully automate the Python library migration process. We systematically approached this goal by first conducting an empirical study to understand the capabilities of LLMs for automating Python library migration. Overall, we find that LLMs can handle complex code changes but we also identify some challenges that need to be addressed to make LLM-based migration practical. Based on these findings and insights, we built MigrateLib, a practical CLI tool that combines LLMs with static and dynamic program analysis to automate the full migration process. Our evaluation on 717 simulated migrations across 175 diverse projects and 48 library pairs shows that MigrateLib can fully automate a large portion of migrations (32%), with most remaining cases requiring only minor manual intervention (median 27% changes). The tool’s layered approach, including post-processing steps, improves LLM migration correctness (from 65% to 86%) and reduces developer effort. Overall, MigrateLib lowers the barrier for large-scale library migration in real-world projects, demonstrating accessible and effective automated library migration.
@@ Acknowledgements.
We thank Dr. Ildar Akhmetov, Northeastern University, for his contributions to the early ideas of this work.
Declarations
Funding
This work was supported by the Canada Research Chairs Program and NSERC CREATE.
Author contributions
Below are the contributions of each author –
  • Mohayeminul Islam contributed to idea conceptualization, design and implementation of the MigrateLib, conducting the experiments, writing script for the experiments, manual evaluations in Phase 1, manual migration fix in Phase 3, and writing the manuscript.
  • Ajay Kumar Jha contributed to idea conceptualization, design of the MigrateLib, manual evaluations in Phase 1, and writing the manuscript.
  • May Mahmoud manual evaluations in Phase 1, and revising the manuscript.
  • Sarah Nadi contributed to idea conceptualization, design of the MigrateLib, manual evaluations in Phase 1, and writing the manuscript. She is the Principal Investigator of the project.
Data availability
The source code of MigrateLib, the experimental data, and the detailed results of our experiments are available at the following repository: https://figshare.com/account/articles/29602517.
Conflict of interest
No conflict.
Ethical approval/informed concent/clinical trial number
Not applicable.